OpenStudy (anonymous):

Which is a factor of 15xy - 45x - 6y + 18?

OpenStudy (anonymous):

y - 2 5x - 3 5x - 2 y - 6

OpenStudy (anonymous):

@cp9454

OpenStudy (anonymous):

Try to factor by grouping: Rewrite 15xy-45x -6y + 18 = (15xy-45x) + (-6y+18) = ... Can you take things from here ?

OpenStudy (anonymous):

honestly I have no idea... @christopherToni

OpenStudy (anonymous):

but why is that the answer

OpenStudy (anonymous):

Okay, focusing on how I grouped things, 15x is a common factor of the first set of parenthesis, and -6 is a common factor of the second set of parenthesis. With this, we see that (15xy-45x) + (-6y+18) = 15x(y-3) -6(y-3) = (15x-6)(y-3) = 3(5x-2)(y-3). So which of these factors appear as one of your choices? :-)

OpenStudy (anonymous):

5x - 2

OpenStudy (anonymous):

Yup, that's correct.